# Entanglement cones and horizons in analogue cosmological production of Dirac fermions

### v1

@@ Abstract
Although gravitational particle production in curved quantum field theories (cQFTs) is key to our understanding of the early universe and black hole physics, its direct observation requires extreme conditions or unrealistic sensitivities. Recent progress in quantum simulators indicates that analogues of gravitational particle production can be observed in table-top experiments of cold atomic gases described by effective cQFTs. This promises a high degree of tunability in the synthesised curved spacetimes and, moreover, sets a clear roadmap to explore the interplay of gravitational particle production with other non-perturbative effects genuine to interacting QFTs. We hereby focus on the appearance of fermion condensates for self-interacting Dirac fermions, and study how dynamical mass generation and spontaneous symmetry breaking affects real-time dynamics. We use the entanglement contour as a tool to analyze the spatio-temporal causal structure of particle production, showing how it can account for the cosmological horizon in accelerating spacetimes, while also being sensitive to the effect of different symmetry-breaking processes. In particular, we show that the combined breakdown of time-reversal symmetry due to the expanding spacetime, and parity due to a pseudo-scalar condensate, manifest through the structure of the light-cone-like propagation of entanglement in this analogue cQFT.
preprint: APS/123-QED
@@
Introduction.– Quantum simulators (QSs) promise to become a transformative tool to explore the complex quantum many-body frontier common to various disciplines of contemporary science  [1, 2, 3]. In theoretical physics, QSs can sometimes realize analogues of phenomena that would otherwise be difficult to observe, if not altogether impossible, in their original realm. These two aspects come together in the study of quantum field theories (QFTs) in curved spacetimes [4, 5, 6, 7, 8, 9, 10] by means of experimental quantum devices [11, 12, 13, 14, 15, 16] that can be used to explore foundational questions about gravity, or test our current cosmological models of the universe. These QSs foster a genuinely multidisciplinary approach that integrates concepts from atomic physics, quantum optics and quantum information theory, with those of high-energy physics and general relativity. A particularly-relevant example is that of quantum entanglement [17, 18, 19], which provides insights into the role of quantum correlations in the complexity frontier of many-body systems, and plays an important role in questions as diverse as the nature of the QFT vacuum or the evaporation of black holes, and also possibly in the information paradox for the structure of spacetime at the quantum level.
@@
In this Letter, we focus on QSs of ultra-cold atoms in Raman optical lattices [20, 21, 22, 23, 24, 25, 26, 27] as analogues of particle production of self-interacting Dirac fermions in an effective spacetime with a tunable accelerated expansion. We demonstrate how the entanglement contour [28], when combined with a non-perturbative variational ansatz based on fermionic Gaussian states [29, 30], can provide a clean spatio-temporal resolution for the real-time dynamics of these Dirac fields. Looking at particle production through this entanglement-contour lens, we unveil interesting phenomena like a cosmological particle horizon in the entanglement spreading, and the impact on the entanglement-cone structure of symmetry-breaking fermion condensates during particle production, pointing at future research opportunities in analogue gravity [31].
@@
The model. We consider a Friedmann-Robertson-Walker spacetime $\mathord{\rm d}s^{2}=\d{d}t^{2}-\mathsf{a}^{2}(t)\mathord{\rm d}\Sigma^{2}$ with a scale factor $\mathsf{a}(t)$. In a cosmological scenario [32], the scale factor connects an inflationary epoch to our current dark-energy and dark-matter dominated universe in $D=3+1$ dimensions. In contrast, in an analogue setup, we are free from these constraints, and can: (i) connect two asymptotic Minkowski regions $\mathsf{a}_{0}\to\sf{a}_{f}$ by an exponential scale factor controlled by the Hubble rate $\mathsf{H}$, such that the duration is $\Delta t=\frac{1}{\mathsf{H}}\log\frac{\mathsf{a}_{\rm f}}{\mathsf{a}_{0}}$; (ii) vary the Hubble rate at will, as it is not fixed by cosmological observations, interpolating between adiabatic $\sf{H}\rightarrow 0$ and sudden $\sf{H}\rightarrow\infty$ expansions; and (iii) explore reduced dimensionalities that may lead to simpler analogue experiments. This flexibility extends to the (iv) dynamical quantum fields that fill in this expanding spacetime which, in the standard cosmological model of inflation, correspond to a scalar field under certain self-interaction potentials. For the reasons exposed below, we shall be interested instead in self-interacting Dirac spinors $\hat{\psi}(x)\to\hat{\psi}_{i}(t)$ discretized on a co-moving chain ${x}_{i}=ia$, where $a$ is the lattice spacing and $i\in\mathbb{Z}_{N_{\rm s}}$. By using conformal time via $\eta=\int{\rm d}t/\mathsf{a}(t)$, and rescaling the fields as $\hat{\psi}_{i}\rightarrow\sqrt{\sf{a}(\eta)}\hat{\psi}_{i}$, there is not need to simulate an inflating lattice with increasing proper distance $d_{ij}(t)=\mathsf{a}(t)|x_{i}-x_{j}|$, and the regularised action reduces to a fermionic QFT with explicit dependence on conformal time. Finally, we can also (iv) explore the subsequent evolution of the produced particles within a flat spacetime $\mathsf{a}_{\rm f}$, which will allow us to derive a neat picture of thermalization as these particles fill in the universe.
We work with a Hamiltonian lattice field theory with rescaled spinors and conformal time associated to the original cQFT action that includes vierbein fields and the spin connection [33]. In particular, we employ a Wilson-type regularization to deal with fermion doubling [34] which, in momentum space, yields a free part $\hat{H}=\sum_{\textrm{k}\in\rm BZ}\hat{\psi}_{\textrm{k}}^{\dagger}h_{\textrm{k}}\big{(}m\mathsf{a}(\eta)\big{)}\hat{\psi}_{\textrm{k}}$ with
@@
where the matrices are $\gamma^{0}=\sigma^{z}$, $\gamma^{1}={\rm i}\sigma^{y}$ and $\gamma^{5}=\gamma^{0}\gamma^{1}$. This free field theory has a simple interpretation as a tight-binding model on a synthetic ladder, and can be realised with minor modifications of recent experiments with alkaline atoms in optical Raman lattices [27]. Here, the ladder legs represent the spinor components, encoded in two hyperfine levels of ${}^{87}{\rm Sr}$, and an energy imbalance that accounts for the time-dependent mass $m\mathsf{a}(\eta)$ can be controlled by the Raman lasers detuning.
Being quadratic, Wick’s theorem holds for this part, and the real-time dynamics can be fully characterised through the correlation matrix $\Gamma_{ij}(\eta)=\bra{\psi(\eta)}\hat{\psi}^{\phantom{\dagger}}_{i}\hat{\psi}_{j}^{\dagger}\ket{\psi(\eta)}$. Regarding the full many-body Hilbert space, the dynamics is constrained to a reduced manifold of so-called fermionic Gaussian states (fGS) {|$\{\ket{\psi(\Gamma)}:\,\Gamma\in\textrm{GL}_{4N_{\rm s}\!}(\mathbb{C})\}$ during the expansion. This can then be used as a variational family to approximate genuinely-interacting QFTs. In this limit, particle production [4, 5] can be formalised by a Bogoliubov transformation [35, 36] that acts by mapping |$\ket{\psi(\Gamma_{0})}\mapsto\ket{\psi(\Gamma_{\rm f})}$ producing particles. Taking advantage of the above points (i)-(iv), we can actually go beyond this simple limit and explore non-perturbative effects by including self-interactions according to the Gross-Neveu model [37], namely
$\hat{{V}}_{\rm int}=\frac{g_{0}^{2}}{2}\sum_{i}\left(\hat{\psi}_{i}^{\dagger}\gamma^{0}\hat{\psi}^{\phantom{\dagger}}_{i}\right)^{\!\!2}\!\!,$ (2)
where $g_{0}^{2}$ is the dimensionless coupling strength which, in contrast to to the bare mass, does not get renormalised by the scale factor. In a previous work [38], we have indeed shown that the fGS ansatz can be used to approximate the vacuum |$\ket{\psi(\Gamma_{0})}$, and that it coincides exactly with large-$N$ predictions [39]. The fGS predictions account for the onset of scalar and pseudo-scalar fermion condensates, which are specific elements of the correlation matrix of fermion bilinears $\Sigma={g_{0}^{2}}\sum_{\textrm{k}}\langle\hat{\psi}_{\textrm{k}}^{\dagger}\gamma^{0}\hat{\psi}_{\textrm{k}}\rangle/{2aN_{\rm s}}$, $\Pi={{\rm i}g_{0}^{2}}\sum_{\textrm{k}}\langle\hat{\psi}_{\textrm{k}}^{\dagger}\gamma^{1}\hat{\psi}_{\textrm{k}}\rangle/{2aN_{\rm s}}$. During the expansion, these condensates will evolve non-perturbatively affecting the particle production and vice versa, which we can approximate variationally via our fGS, leading to a set of self-consistent non-linear differential equations
@@
${\rm i}\frac{\mathord{\rm d}}{\mathord{\rm d}\eta}\Gamma(\eta)=\left[\tilde{h}_{\textrm{k}}\big{(}m\mathsf{a}(\eta),\Gamma(\eta)\big{)},\;\Gamma(\eta)\right],$ (3)
where $h_{\textrm{k}}\to\tilde{h}_{\textrm{k}}\big{(}m\mathsf{a}(\eta),\Gamma\big{)}={h}_{\textrm{k}}\big{(}m\mathsf{a}(\eta)+\Sigma(\Gamma)\big{)}-\mathord{\rm i}\Pi(\Gamma)\gamma^{1}.$ We see how the non-linearities arise through $\tilde{h}_{\rm k}$ via the self-energy expressed in terms of the condensates, which follows from Wick’s theorem on the fGS manifold $\bra{\psi(\Gamma)}\hat{{V}}_{\rm int}\ket{\psi(\Gamma)}$.
Figure 1: (a) Spatio-temporal propagation of EC in the asymptotic flat region $\mathsf{a}_{\rm f}$, both in the free $g_{0}^{2}=0$ and in the interacting $g_{0}^{2}=1$ regimes, and setting $ma=-1$, $\mathsf{a}_{0}=0.7$, $\mathsf{a}_{\rm f}=1.3$, ${\sf H}a=100$ (sudden expansion). The dashed white lines indicate the light-cone-like predictions based on the group velocity after the sudden expansion. (b,c) Spatio-temporal propagation of EC during a de Sitter expansion, using conformal and cosmological time respectively, and setting $ma=1$, ${\sf H}a=0.1$, $g_{0}^{2}=2$, $\mathsf{a}_{0}=1/3$, with $\eta$ and $t$ expressed in lattice units. The expansion in (b) starts at $\eta_{0}a=-30$ and extends to the distant future $\eta\to 0^{-}$, approximated by $\eta a=-0.001362$. The shaded regions denote the asymptotic spatial boundary that particles cannot cross. In (c) the corresponding expansion covers $t_{0}=0$ to $ta\to\infty$, here approximated by $t=100/a$.
Solving these equations with the initial condition |$\ket{\psi(\Gamma_{0})}$ one obtains a time-evolved fGS |$\ket{\psi(\Gamma(\eta_{\rm f}))}$ with a non-vanishing density of particles with respect to the instantaneous groundstate |$\ket{\psi(\Gamma_{\rm f})}$ at $(m\mathsf{a}_{\rm f},g_{0}^{2})$. Within our variational approximation, the interpretation in terms of a Bogoliubov transformation is still useful, but it now depends on the full non-perturbative time history $\{\Gamma\}_{\eta}=\{\Gamma(\eta):\,\eta\in[\eta_{0},\eta_{\rm f}]\}$, allowing to express the evolved state as
@@
with normalised Bogoliubov parameters $|\alpha_{\textrm{k}}|^{2}+|\beta_{\textrm{k}}|^{2}=1$. The density of produced fermions after the expansion is $n_{a}=n_{b}=\sum_{\textrm{k}\in{\rm BZ}}|\beta_{\textrm{k}}(\{\Gamma\}_{\eta})|^{2}/{{aN_{\rm s}\mathsf{a}_{\rm f}}}$, which is the result of a back-to-back production of particle-antiparticle pairs. We note that previous works [40, 41, 42, 43] on free QFTs have exploited the formal analogy of this transformation with two-mode squeezing to characterise mode entanglement in a cosmological scenario in momentum space, here connected to entanglement between right- and left-moving excitations. This, however, does not provide any spatial resolution that links to the causal propagation of entanglement, or to the existence of cosmological horizons due to an accelerated expansion. Moreover, to our knowledge, there has been no prior study of the effect of interactions and fermion condensates on the entanglement during gravitational expansion. We fill in both gaps in this work.
Correlation light cones.– Let us start moving away from the aforementioned mode entanglement by focusing on the block entanglement entropy (EE) $S_{A}=-\mathop{\rm Tr}(\rho_{A}\log\rho_{A})$, which quantifies the amount of quantum correlations across the $AB$ bipartition of the system $\rho_{A}=\mathop{\rm Tr}_{B}\ket{\psi}\!\!\bra{\psi}$ |, where $A$ would be a block of $\ell_{A}$ sites. We note that the EE has been studied in great detail for quantum quenches in many-body models, where a system evolves under a sudden change of the Hamiltonian, yielding a neat quasi-particle picture in which a linear entanglement growth is a manifestation of a light-cone-like propagation of excitations [44, 45, 46, 47, 48, 49, 50].

### v2

# A quasi-particle picture for entanglement cones and horizons in analogue cosmology

@@ Abstract
Although particle production in curved quantum field theories (cQFTs) is key to our understanding of the early universe and black hole physics, its direct observation requires extreme conditions or unrealistic sensitivities. Recent progress in quantum simulators indicates that analogues of cosmological particle production can be observed in table-top experiments of cold atomic gases described by effective cQFTs. This promises a high degree of tunability in the synthesised curved spacetimes and, moreover, sets a clear roadmap to explore the interplay of particle production with other non-perturbative effects genuine to interacting QFTs. We hereby focus on the appearance of scalar and pseudo-scalar condensates for self-interacting Dirac fermions, and study how dynamical mass generation and spontaneous symmetry breaking affects real-time dynamics through the lens of entanglement. We use the entanglement contour (EC) to analyse the spatio-temporal structure of particle production, showing that a quasi-particle picture for the EC captures the cosmological horizon in accelerating spacetimes, while also being sensitive to the effect of different symmetry-breaking processes. In particular, we show that the combined breakdown of time-reversal symmetry due to the expanding spacetime, and parity due to a pseudo-scalar condensate, manifests through the structure of the light-cone-like propagation of entanglement.
@@
Introduction.– Quantum simulators (QSs) promise to become a transformative tool to explore the complex quantum many-body frontier common to various disciplines of contemporary science [1, 2, 3]. In theoretical physics, QSs can sometimes realise analogues of phenomena that would otherwise be difficult to observe in their original realm, if not altogether impossible. These two aspects come together in the study of quantum field theories (QFTs) in curved spacetimes [4, 5, 6, 7, 8, 9, 10] by means of experimental quantum devices [11, 12, 13, 14, 15, 16] that can be used to explore foundational questions about gravity, or test our current cosmological models of the universe. These QSs foster a genuinely multidisciplinary approach that integrates concepts from atomic physics, quantum optics and quantum information theory, with those of high-energy physics and general relativity. A particularly relevant example is that of quantum entanglement [17, 18, 19], which provides insights into the role of quantum correlations in the complexity frontier of many-body systems, and plays an important role in questions as diverse as the nature of the QFT vacuum or the evaporation of black holes, and also possibly in the information paradox for the structure of spacetime at the quantum level.
@@
In this Letter, we focus on QSs of ultra-cold atoms in Raman optical lattices [20, 21, 22, 23, 24, 25, 26, 27] as analogues of particle production of self-interacting Dirac fermions in an expanding spacetime with a tunable scale factor. We demonstrate how the entanglement contour [28], when combined with the quasi-particle picture of entanglement dynamics [29, 30, 31, 32] and a non-perturbative variational ansatz based on fermionic Gaussian states [33, 34], can provide a clean spatio-temporal resolution for the real-time dynamics. Looking at particle production through the entanglement contour, we unveil interesting phenomena like a cosmological particle horizon in the entanglement spreading, and the impact of symmetry-breaking fermion condensates during particle production on the entanglement-cone structure, pointing at future research opportunities in analogue gravity [35].
@@
The model. We consider a Friedmann-Robertson-Walker spacetime $\mathord{\rm d}s^{2}=\mathord{\rm d}t^{2}-\mathsf{a}^{2}(t)\mathord{\rm d}\Sigma^{2}$ with a scale factor $\mathsf{a}(t)$ and $\mathord{\rm d}\Sigma^{2}$ depends on the metric of the spatial slice. In a cosmological scenario [36], the scale factor connects an inflationary epoch to our current dark-energy and dark-matter dominated universe in $D=3+1$ dimensions. In contrast, in an analogue setup, we are free from these constraints, and can: (i) connect two asymptotic Minkowski regions $\mathsf{a}_{0}\to\sf{a}_{\rm f}$ by an exponential scale factor controlled by the Hubble rate $\mathsf{H}$, such that the duration is $\Delta t=\frac{1}{\mathsf{H}}\log\frac{\mathsf{a}_{\rm f}}{\mathsf{a}_{0}}$; (ii) vary the Hubble rate at will, as it is not fixed by cosmological observations, interpolating between adiabatic $\sf{H}\rightarrow 0$ and sudden $\sf{H}\rightarrow\infty$ expansions; and (iii) explore reduced dimensionalities that may lead to simpler analogue experiments, such as $\mathord{\rm d}\Sigma^{2}=\mathord{\rm d}{\rm x}^{2}$ in $D=1+1$ dimensions. This flexibility extends to the dynamical quantum fields evolving in this expanding spacetime, which, in the standard cosmological model of inflation, is driven by a scalar field governed by certain self-interaction potentials. For the reasons exposed below, we shall be interested in self-interacting Dirac spinors ${\psi}(x)\to{\psi}_{i}(t)$ discretised on a co-moving chain ${x}_{i}=ia$, where $a$ is the lattice spacing and $i\in\mathbb{Z}_{N_{S}}$. By using conformal time via $\eta=\int{\rm d}t/\mathsf{a}(t)$, and rescaling the fields as $\psi_{i}\rightarrow\sqrt{\sf{a}(\eta)}\psi_{i}$, there is no need to simulate an inflating lattice with increasing proper distance $d_{ij}(t)=\mathsf{a}(t)|x_{i}-x_{j}|$, and the regularised action reduces to a fermionic QFT with explicit dependence of the microscopic parameters on conformal time. Finally, we can also (iv) explore the subsequent evolution of the produced particles within a flat spacetime $\mathsf{a}_{\rm f}$, which will allow us to derive a neat picture of thermalisation as these particles propagate and fill in the universe.
We work with a Hamiltonian lattice field theory that can be derived from the original cQFT action
$S=\!\int\!{\rm d}^{D}x\sqrt{-g}\,\overline{\psi}(x)\Big(g_{\mu\nu}(x)\tilde{\gamma}^{\mu}\!(x)\nabla_{\mu}-m\!\Big)\!\psi(x)-V_{\rm int}(\overline{\psi},{\psi}),$ (1)
which includes a generic metric $g_{\mu\nu}(x)$ with a determinant $g={\rm det}(g_{\mu\nu})$ setting the volume element, curved gamma matrices fulfilling $\{\tilde{\gamma}^{\mu}\!(x),\tilde{\gamma}^{\nu}\!(x)\}=2g^{\mu\nu}(x)$, and a covariant derivative $\nabla_{\mu}$ that contains the spin connection, and the Christoffel symbols [37]. We also include fermion self-interactions through a potential $V_{\rm int}$ that depends on the fermion field and its adjoint $\overline{\psi}(x)=\psi^{\dagger}(x)\gamma^{0}$. Working in conformal time and with the above rescaled spinors, the Hamiltonian adopts a very transparent form which can then be regularised in the above co-moving lattice. In particular, we employ a Wilson-type regularisation to deal with fermion doubling [38] which, in momentum space, yields a free part ${H}=\sum_{\textrm{k}\in\rm BZ}\psi_{\textrm{k}}^{\dagger}h_{\textrm{k}}\big(m\mathsf{a}(\eta)\big)\psi_{\textrm{k}}$ with
@@
where the gamma matrices are $\gamma^{0}=\sigma^{z}$, $\gamma^{1}=\mathord{\rm i}\sigma^{y}$ and $\gamma^{5}=\gamma^{0}\gamma^{1}$. This free field theory has a simple interpretation as a tight-binding model on a synthetic ladder, and can be realised with minor modifications of recent experiments with alkaline-earth atoms in Raman lattices [27]. Here, the ladder legs represent the spinor components, encoded in two hyperfine levels of $^{87}{\rm Sr}$, and an energy imbalance that accounts for the time-dependent mass $m\mathsf{a}(\eta)$, controlled by the Raman detuning.
[…]
${{V}}_{\rm int}=\frac{g_{0}^{2}}{2}\sum_{i}\left(\psi_{i}^{\dagger}\gamma^{0}\psi^{\phantom{\dagger}}_{i}\right)^{\!\!2}\!\!.$ (3)
[…]
$\Sigma(\eta)=\frac{g_{0}^{2}}{2aN_{S}}\sum_{\textrm{k}}\langle\psi_{\textrm{k}}^{\dagger}\gamma^{0}\psi_{\textrm{k}}\rangle,\hskip 6.45831pt\Pi(\eta)=\frac{\mathord{\rm i}g_{0}^{2}}{2aN_{S}}\sum_{\textrm{k}}\langle\psi_{\textrm{k}}^{\dagger}\gamma^{1}\psi_{\textrm{k}}\rangle.$ (4)
During the expansion, these condensates will evolve non-perturbatively affecting the particle production and vice versa, which we can approximate variationally via our fGS, leading to a set of self-consistent non-linear differential equations
@@
$\mathord{\rm i}\frac{\mathord{\rm d}}{\mathord{\rm d}\eta}\Gamma(\eta)=\left[\tilde{h}_{\textrm{k}}\big(m\mathsf{a}(\eta),\Gamma(\eta)\big),\;\Gamma(\eta)\right].$ (5)
[…]
@@
with normalised Bogoliubov parameters $|\alpha_{\textrm{k}}|^{2}+|\beta_{\textrm{k}}|^{2}=1$. The density of produced fermions after the expansion is $n_{a}=n_{b}=\sum_{\textrm{k}\in{\rm BZ}}|\beta_{\textrm{k}}(\{\Gamma\}_{\eta})|^{2}/{{aN_{S}\mathsf{a}_{\rm f}}}$, which is the result of a back-to-back production of particle-antiparticle pairs. We note that previous works [44, 45, 46, 47] on free QFTs have exploited the formal analogy of this transformation with two-mode squeezing to characterise mode entanglement in a cosmological scenario in momentum space, capturing in this way correlations between right- and left-moving excitations. This, however, does not provide any spatial resolution that links to the causal propagation of entanglement, or to the existence of cosmological horizons due to an accelerated expansion. Moreover, to our knowledge, there has been no prior study of the effect of interactions and fermion condensates on the entanglement during gravitational expansion. We fill in both gaps in this work.
Entanglement and the quasi-particle picture.– Let us start moving away from the aforementioned mode entanglement by focusing on the block entanglement entropy (EE) $S_{A}=-\mathop{\rm Tr}(\rho_{A}\log\rho_{A})$, which quantifies the amount of quantum correlations across the $AB$ bipartition of the system $|\rho_{A}=\mathop{\rm Tr}_{B}\ket{\psi}\!\!\bra{\psi}$, where $A$ would be a block of $\ell_{A}$ sites. We note that the EE has been studied in great detail under the so-called quasi-particle picture. In its original realm, the quasi-particle picture describes the evolution of entanglement entropy as carried by propagating quasi-particles produced in pairs after a quantum quench. It was first introduced as a qualitative explanation for homogeneous quantum quenches in quantum chains [29], and then quantitatively demonstrated for the XY model [30] and other generic integrable models [31, 32]. It has also been studied for other entanglement measurements beyond entanglement entropy, such as the so-called entanglement links [48]. This picture suggests that, after a quantum quench, the initial state acts as a reservoir of excited quasi-particles that are created in pairs and propagate in straight lines with opposite velocities given by $v_{\textrm{k}}=\pm|\partial_{\textrm{k}}\varepsilon_{\textrm{k}}|$, where $\varepsilon_{\textrm{k}}$ is an energy dispersion relation for the quasi-particle labelled by k. From this perspective, when the partition involves two spatial regions $A$ and $B$, entanglement growth is due to entangled pairs of propagating quasi-particles residing in different sides of the partition. In our context of conformal time, the quasi-particle picture of the EE dynamics is
$S_{A}(\eta)=\eta\int_{v_{\textrm{k}}>\frac{\ell_{A}}{2\eta}}\frac{\mathord{\rm d}\textrm{k}}{2\pi}\,2v_{\textrm{k}}s(\textrm{k})+\ell_{A}\int_{v_{\textrm{k}}<\frac{\ell_{A}}{2\eta}}\frac{\mathord{\rm d}\textrm{k}}{2\pi}\,s(\textrm{k}),$ (7)
where the first term describes an initial linear increase of entanglement over conformal time, which accounts for a phase in which quasi-particles are entering in the $A$ block. The second term, on the other hand, describes an equilibration to an extensive amount of entanglement proportional to $\ell_{A}$ that appears for sufficiently long times, such that the same amount of quasi-particles enter as those that exit the block. Note that, if there is a maximum propagation velocity, the limits of the integrals define an effective causal cone for the spread of entanglement in the system. In this formula, $s(\textrm{k})$ denotes the amount of entanglement carried by the pair of quasi-particles, and can be obtained by analysing the long-time behaviour of the quantum state, which is expected to approach a generalised Gibbs ensemble (GGE) [49] with a well known thermodynamic entropy. Alternatively, this quantity can be calculated exactly for some integrable systems by knowing the specific form of an initial squeezed state [50].
Let us now adapt this quasi-particle picture to our cosmological scenario, noting that the Hubble rate $\mathsf{H}$ allows us to smoothly interpolate between the adiabatic and quench limits. We note that after a antiparticle-hole transformation [51], quasi-particles correspond to pairs of fermions and antifermions created back-to-back with opposite quasi-momentum, as described by Eq. (6). For this case, closed formulas can be found for the mode entanglement between particles and antiparticles, which are derived by tracing out the sector of antiparticles from the density matrix. We have
$\rho=\bigotimes_{\textrm{k}}\rho_{\textrm{k}}=\bigotimes_{\textrm{k}}\bigg(|\alpha_{\textrm{k}}|^{2}|0\rangle\langle 0|+|\beta_{\textrm{k}}|^{2}|1_{\textrm{k}},\bar{1}_{-\textrm{k}}\rangle\langle 1_{\textrm{k}},\bar{1}_{-\textrm{k}}|\bigg),$ (8)
where we have defined $|1_{\textrm{k}},\bar{1}_{-\textrm{k}}\rangle={a}^{\dagger}_{\textrm{k}}{b}^{\dagger}_{-\textrm{k}}|0\rangle$. Then, tracing out the antiparticles sector, we arrive at
$\rho^{\rm part}_{\textrm{k}}=|\alpha_{\textrm{k}}|^{2}|0\rangle\langle 0|+|\beta_{\textrm{k}}|^{2}|1_{\textrm{k}}\rangle\langle 1_{\textrm{k}}|.$ (9)
The entanglement entropy between particles and antiparticles in the mode k is therefore given by the von Neumann entropy of $\rho^{\rm part}_{\textrm{k}}$, $S(\rho_{\textrm{k}}^{\rm part})=-\mathop{\rm Tr}(\rho_{\textrm{k}}^{\rm part}\log\rho_{\textrm{k}}^{\rm part})$, which in terms of the Bogoliubov coefficients is given by
$S(\rho_{\textrm{k}}^{\rm part})=-|\alpha_{\textrm{k}}|^{2}\log|\alpha_{\textrm{k}}|^{2}-|\beta_{\textrm{k}}|^{2}\log|\beta_{\textrm{k}}|^{2}.$ (10)
The relation between the Bogoliubov parameters for fermions, $|\alpha_{\textrm{k}}|^{2}+|\beta_{\textrm{k}}|^{2}=1$, allows to construct this expression using only the spectrum of particle production, with entanglement entropy being upper-bounded by $S(\rho^{\rm part})\leq\log 2$ for $|\beta_{\textrm{k}}|^{2}=\textstyle\frac{1}{2}$. Using this expression, the contributions $s(\textrm{k})$ in Eq. (7) are found to be $s(\textrm{k})=2S(\rho^{\rm part})$, where the factor of 2 takes into account the two possibilities of a particle propagating out of the partition while an antiparticle stays inside and vice versa.
This quasi-particle prediction (7) becomes exact in the non-interacting regime $g_{0}^{2}=0$ and in the scaling limit $\ell_{A},\;t\to\infty$, while keeping $\ell_{A}/t$ constant. For large but finite $t$ and $\ell_{A}$, it gives the leading order contribution for the EE dynamics of our propagating Dirac fields. In Fig. 1 we benchmark the quasi-particle prediction with the exact time dynamics of the of the EE for two different values of $\ell_{A}$. We observe some deviations for small times, but a good agreement with Eq. (7) for increasingly larger partition size as one approaches the scaling limit, according to the expected result.
Figure 1: Time evolution of entanglement entropy $S_{A}(\eta)$ for a partition of (a) $\ell_{A}=32$ sites from a $N_{S}=128$ sites system, and (b) $\ell_{A}=128$ sites from a $N_{S}=512$ sites system. The blue line represents the numerical result while the red dashed line represents the quasi-particle prediction. The parameters are set as $ma=1$, $g_{0}^{2}=0$, $\mathsf{a}_{0}=0.01$ and $\mathsf{a}_{\rm f}=10$.
Entanglement contour and causal cones.– One limitation of the block EE is the lack of spatial resolution even in the quasi-particle regime. This resolution would be desirable as it would yield information about the real-time propagation of pairs of particle-antiparticle, giving insights to the causal structure of the expanding spacetime. This limitation can be circumvented by resorting to the concept of entanglement contour (EC) [28], which assigns to each site $i$ of the partition a specific contribution $S_{i}$ to the overall EE, such that $\sum_{i}S_{i}=S_{A}$. With it, one can get further information on the local distribution of entanglement within the block, allowing us to bring this quasi-particle picture to a new domain. In particular, we show below that the EC allows us to identify a cosmological particle horizon due to the accelerating spacetime.
Although a general recipe for the calculation of EC in arbitrary interacting models is lacking, the fGS ansatz provides a well-defined approximation to the EC via the aforementioned correlation matrix $\Gamma(\eta)$, as the spectrum of the block density matrix $\rho_{A}$ is determined by the restricted correlation matrix $(\Gamma_{A}(\eta))_{ij}=\Gamma_{ij}(\eta)|_{i,j\in A}$ [52, 53]. Diagonalising this matrix by a unitary transformation $U^{\dagger}\Gamma_{A}U={\rm diag}(\{\nu_{\textrm{k}}\}_{\textrm{k}\in{\rm BZ}})$, one obtains an EC $S_{i}(\eta)=\sum_{\textrm{k}\in{\rm BZ}}p_{i}(\textrm{{k}},\eta)s_{\textrm{k}}(\eta)$ that weighs each contribution $s_{\textrm{k}}(\eta)=-\nu_{\textrm{k}}(\eta)\log\nu_{\textrm{k}}(\eta)-(1-\nu_{\textrm{k}}(\eta))\log(1-\nu_{\textrm{k}}(\eta))$ at the $i$-th site by $p_{i}(\textrm{k},\eta)=|U_{\textrm{k}i}(\eta)|^{2}$. Since this matrix is determined self-consistently by solving for the fGS differential equations (5), we can explore how entanglement evolves in the spacetime including interactions beyond perturbation theory, e.g. symmetry-breaking fermion condensates.
The quasi-particle picture can be extended to the entanglement contour, yielding the leading order contribution to the numerical results in the scaling limit. To do this, one starts by stating that pairs of particle-antiparticle are created homogeneously along the partition. Those created at sites closer to the edges will contribute earlier to the increase of entanglement, and effectively one sees entanglement stemming from both boundaries with a maximum velocity given by $v_{\rm max}=2\max_{\textrm{k}\in{\rm BZ}}v_{\textrm{k}}$. The factor of $2$ simply manifests the fact that both quasi-particles travel oppositely with the same velocity, so entanglement effectively travels at twice the speed of the individual quasi-particles. This behaviour is captured by the following expression
$S_{A}(\mathbf{x})=\int\frac{d\textrm{k}}{2\pi}\,s(\textrm{k})\big(\Theta(2v_{\textrm{k}}\eta-\mathbf{x})+\Theta(2v_{\textrm{k}}\eta-(\ell_{A}-\mathbf{x}))\big),$ (11)
where $\Theta(x)$ is the Heaviside step function. This picture translates naturally to our lattice field theory as, by virtue of the particle-hole symmetry, EC is enforced to be invariant upon the exchange of the spinor components $\alpha\in\{u,\;d\}$, such that Eq. (11) distributes evenly between them, $S_{A}({\bf x},\alpha)=\textstyle\frac{1}{2}S_{A}({\bf x})$.
Figure 2: Spatio-temporal propagation of EC in the asymptotic flat region $\mathsf{a}_{\rm f}$, both in the free $g_{0}^{2}=0$ and in the interacting $g_{0}^{2}=1$ regimes, and setting ma=1ma=-1, $\mathsf{a}_{0}=0.7$, $\mathsf{a}_{\rm f}=1.3$, ${\sf H}a=100$ (sudden expansion). The dashed white lines indicate the light-cone-like predictions based on the group velocity after the sudden expansion.